The difference between the ages of two men is 10 years.15 years ago the older was twice the younger .What are their present ages?
Answer by ankor@dixie-net.com(22740)

let a = the age of the older man
:
"The difference between the ages of two men is 10 years."
a - 10 = age of the younger
:
"15 years ago the older was twice the younger ."
a - 15 = 2(a-10-15)
a - 15 = 2(a-25)
a = 2a - 50 + 15
a = 2a - 35
35 = 2a - a
a = 35 now
:
What are their present ages?